# HRI-RECAPP-2024-03

### v1

@@ I Introduction
One of the earliest proposed BSM scenarios was Supersymmetry (SUSY) - in addition to stabilizing the running of the Higgs mass, SUSY theories also offered the prospect of gauge coupling unification. Considerations of anomaly cancellation necessitated that these theories necessarily invoke two Higgs doublets to engineer electroweak symmetry breaking (EWSB) and provide masses to the various gauge bosons and fermions. Along these lines, models with extended scalar sectors came to be also constructed as these were interesting in their own right. The earliest such attempts were along the lines of the “Two Higgs Doublet Models” (2HDM) [6] - as the name suggests, these have two Higgs doublets and have been studied extensively in the literature [7, 8, 9]. The study of extended Higgs sectors is by no means confined to only the 2HDMs - the Georgi-Machacek model (which includes a Higgs triplet as well) [10, 11, 12], various other avatars of the 2HDM with extra inert scalars to serve as dark matter candidates etc. have also been studied. Interestingly, the Higgs data still leaves room for many of these extended Higgs sector scenarios and these are not (yet) completely ruled out.
Recently, the “Three Higgs Doublet” models (3HDM) have also been introduced. These have the attractive possiblity of allowing for CP violation [13, 14] in addition to offering a rich scope for phenomenology [15, 16] given that they have many as yet undiscovered scalar particles in their spectrum, and dark matter prospects [17, 18, 19]. There are many versions of 3HDM one can construct based on whether one wants EWSB to proceed via all three doublets or just one or two of them. In addition, various possibilities exist for patterns of Yukawa couplings in these models each with their own unique phenomenology. There are many papers in the literature that look at one or more aspects of the 3HDM [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 14], and also address the possibility of whether the parameter space of the 3HDM is allowed with respect to specific theoretical (like unitarity and perturbativity considerations) and experimental constraints. Our goal in this paper is to look at a particular version of the 3HDM and analyze its parameter space, subjecting it to all known theoretical and experimental constraints. As we will lay out in the paper, this is a rather daunting task given the large number of parameters in the model made only tractable by the use of active learning algorithms.
@@ II.4 Couplings
$\bar{t}bH_{2}^{+}$          $-c_{\beta_{2}}s_{\gamma_{2}}y^{u}_{3}P_{L}-(c_{\beta_{1}}c_{\gamma_{2}}+s_{\beta_{1}}s_{\beta_{2}}s_{\gamma_{2}})y^{d*}_{3}P_{R}$
$\bar{t}bH_{3}^{+}$          $c_{\beta_{2}}c_{\gamma_{2}}y^{u}_{3}P_{L}-(c_{\beta_{1}}s_{\gamma_{2}}-s_{\beta_{1}}s_{\beta_{2}}c_{\gamma_{2}})y^{d*}_{3}P_{R}$
@@ III.2 Medial Hierarchy
Again le $k=\cos(\alpha_{1}-\beta_{1})$, this can be re-written as
$-c_{\beta_{2}}c_{\alpha_{3}}\sqrt{1-k^{2}}+s_{\beta_{2}}c_{\alpha_{2}}s_{\alpha_{3}}-c_{\beta_{2}}s_{\alpha_{2}}s_{\alpha_{3}}k=1$ (21)
@@
  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$. It is clear from Eqn. 21 that in this case, we have $\sin\alpha_{3}\,\sin(\beta_{2}-\alpha_{2})=0$. This has two possible solutions: $\alpha_{3}=\frac{\pi}{2}$, $\alpha_{2}=\beta_{2}-\frac{\pi}{2}$ and $\alpha_{3}=\frac{3\pi}{2}$, $\alpha_{2}=\beta_{2}-\frac{3\pi}{2}$. We can combine these conditions to read $\tan\alpha_{2}=-\cot\beta_{2}$.
@@ III.3 Inverted Hierarchy
$c_{\alpha_{3}}s_{\beta_{2}}c_{\alpha_{2}}-c_{\alpha_{3}}c_{\beta_{2}}s_{\alpha_{2}}k+c_{\beta_{2}}s_{\alpha_{3}}\sqrt{1-k^{2}}=1.$ (23)
@@
  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$. It is clear from Eqn. 23, that in this case we have $\cos\alpha_{3}\,\sin(\beta_{2}-\alpha_{2})=0$. This has two possible solutions: $\alpha_{3}=0$, $\alpha_{2}=\beta_{2}-\frac{\pi}{2}$ and $\alpha_{3}=\pi$, $\alpha_{2}=\beta_{2}-\frac{3\pi}{2}$. We can combine these conditions to read $\tan\alpha_{2}=-\cot\beta_{2}$.
@@ V Methodology
The Three Higgs Doublet Model (3HDM) presents a formidable challenge due to the curse of dimensionality, with its scalar potential governed by $14$ independent parameters. Our objective is to efficiently explore this $14$ dimensional parameter space to identify regions that satisfy theoretical constraints. Exhaustively searching for allowed points using a grid-based method would be computationally infeasible, potentially taking years. To circumvent this, we propose training a binary machine learning classifier  [^4] For a comprehensive review of ML techniques in particle physics, the reader is encouraged to consult [45, 46]. to predict whether a given point adheres to the constraints. However, the initial dataset generation reveals a substantial class imbalance, where most points fall into the disallowed category, thus skewing the dataset. The success of the machine learning model hinges on both the quality and representativeness of the training data. Crucially, points near the decision boundary—where the classifier exhibits the greatest uncertainty—carry more informational value than others. To optimize the selection of such data points, we employ an Active Learning [47] strategy. Our Active Learning algorithm significantly enhances the efficiency of the neural network classifier by focusing on a minimal yet highly informative subset of points. This process unfolds as follows:
@@ VII Discussion and Conclusions
To understand the regions of parameter space that are allowed by all experimental and theoretical constraints, we studied three cases: Alignment-1 wherein the lightest of the three Higgses is the SM-like Higgs, Alignment-2 with the second lightest being the 125 GeV Higgs, and Alignment-3 in which the heaviest of the three is identified as the SM-like Higgs. Imposing all theoretical and experimental and theoretical constraints even in this simplified set-up is still a daunting task given the multidimensional nature of the 3HDM parameter space. To mitigate this, we adopt a machine learning strategy by designing an active learning algorithm to effectively delineate the decision boundary between allowed and disallowed regions. This reduces the computational time significantly, while also maximing accuracy.
We find, under this set-up, that this model - though highly constrained, still does have a lot of parameter space that can potentially be probed at the LHC and other future colliders. Both the Alignment-1 and Alignment-2 cases have heavy scalar and pseudoscalar Higgs bosons in the range of $100-600$ GeV, potentially in the right ball park to warrant a dedicated LHC study. Interestingly, the case of having at least one CP-even Higgs lighter than the SM-like 125 GeV Higgs is not ruled out, as our results for the Alignment-2 case demonstrate. However, the more radical scenario of having two Higgses lighter than the 125 GeV SM-like Higgs is not favored by electroweak precision data. Nevertheless, the 3HDM remains a very interesting possibility in its own right, both from the point of view of the theoretical formulation (that can take many avatars depending on the discrete symmetries and the choice of Yukawas) and the associated collider phenomenology.
@@ Acknowledgements.
A.K. acknowledges the support from the Director’s Fellowship at IIT Gandhinagar. A.S. thanks Anusandhan National Research Foundation (ANRF) for providing necessary financial support through the SERB-NPDF grant (Ref No: PDF/2023/002572). S.K.R. acknowledges the support from the Department of Atomic Energy (DAE), India, for the Regional Centre for Accelerator-based Particle Physics (RECAPP), Harish Chandra Research Institute.
@@ Appendix A Charged Higgs Mass Matrix Calculation
$V_{C}^{mass}\supset\begin{pmatrix}\phi_{1}^{-}&\phi_{2}^{-}&\phi_{3}^{-}\end{pmatrix}\frac{\mathcal{M}^{2}_{\phi^{\pm}}}{2}\begin{pmatrix}\phi_{1}^{+}\\ \phi_{2}^{+}\\ \phi_{3}^{+}\end{pmatrix}$
@@ A.1 Transformation
$V_{C}^{mass}=\begin{pmatrix}\phi_{1}^{-}&\phi_{2}^{-}&\phi_{3}^{-}\end{pmatrix}\frac{\mathcal{M}^{2}_{\phi^{\pm}}}{2}\begin{pmatrix}\phi_{1}^{+}\\ \phi_{2}^{+}\\ \phi_{3}^{+}\end{pmatrix}$

### v2

# HRI-RECAPP-2025-03

@@ I Introduction
One of the earliest proposed BSM scenarios was Supersymmetry (SUSY) - in addition to stabilizing the running of the Higgs mass, SUSY theories also offered the prospect of gauge coupling unification. Considerations of anomaly cancellation necessitated that these theories necessarily invoke two Higgs doublets to engineer electroweak symmetry breaking (EWSB) and provide masses to the various gauge bosons and fermions. Along these lines, models with extended scalar sectors came to be also constructed as these were interesting in their own right. The earliest such attempts were along the lines of the “Two Higgs Doublet Models” (2HDM) [6] - as the name suggests, these have two Higgs doublets and have been studied extensively in the literature [7, 8, 9]. The study of extended Higgs sectors is by no means confined to only the 2HDMs - the Georgi-Machacek model (which includes a Higgs triplet as well) [10, 11, 12], various other avatars of the 2HDM with extra inert scalars to serve as dark matter candidates [13, 14, 15, 16] etc. have also been studied. Interestingly, the Higgs data still leaves room for many of these extended Higgs sector scenarios and these are not (yet) completely ruled out.
Recently, the “Three Higgs Doublet” models (3HDM) have also been introduced. These have the attractive possiblity of allowing for CP violation [17, 18] in addition to offering a rich scope for phenomenology [19, 20] given that they have many as yet undiscovered scalar particles in their spectrum, and dark matter prospects [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. There are many versions of 3HDM one can construct based on whether one wants EWSB to proceed via all three doublets or just one or two of them. In addition, various possibilities exist for patterns of Yukawa couplings in these models each with their own unique phenomenology. There are many papers in the literature that look at one or more aspects of the 3HDM [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 18, 45], and also address the possibility of whether the parameter space of the 3HDM is allowed with respect to specific theoretical (like unitarity and perturbativity considerations) and experimental constraints. Our goal in this paper is to look at a particular version of the 3HDM and analyze its parameter space, subjecting it to all known theoretical and experimental constraints. As we will lay out in the paper, this is a rather daunting task given the large number of parameters in the model made only tractable by the use of active learning algorithms.
@@ II.4 Couplings
$\bar{t}bH_{2}^{+}$          $-c_{\beta_{2}}s_{\gamma_{2}}y^{u}_{3,3}V_{CKM(3,3)}P_{L}-(c_{\beta_{1}}c_{\gamma_{2}}+s_{\beta_{1}}s_{\beta_{2}}s_{\gamma_{2}})y^{d*}_{3,3}V_{CKM(3,3)}P_{R}$
$\bar{t}bH_{3}^{+}$          $c_{\beta_{2}}c_{\gamma_{2}}y^{u}_{3,3}V_{CKM(3,3)}P_{L}-(c_{\beta_{1}}s_{\gamma_{2}}-s_{\beta_{1}}s_{\beta_{2}}c_{\gamma_{2}})y^{d*}_{3,3}V_{CKM(3,3)}P_{R}$
@@ III.2 Medial Hierarchy
Again let $k=\cos(\alpha_{1}-\beta_{1})$, this can be re-written as
±$\pm c_{\beta_{2}}c_{\alpha_{3}}\sqrt{1-k^{2}}+s_{\beta_{2}}c_{\alpha_{2}}s_{\alpha_{3}}-c_{\beta_{2}}s_{\alpha_{2}}s_{\alpha_{3}}k=1$ (21)
@@
  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$. It is clear from Eqn. 21 that in this case, we have $\sin\alpha_{3}\,\sin(\beta_{2}-\alpha_{2})=1$. This has two possible solutions: $\alpha_{3}=\frac{\pi}{2}$, $\alpha_{2}=\beta_{2}-\frac{\pi}{2}$ and $\alpha_{3}=\frac{3\pi}{2}$, $\alpha_{2}=\beta_{2}-\frac{3\pi}{2}$. We can combine these conditions to read $\tan\alpha_{2}=-\cot\beta_{2}$.
@@ III.3 Inverted Hierarchy
$c_{\alpha_{3}}s_{\beta_{2}}c_{\alpha_{2}}-c_{\alpha_{3}}c_{\beta_{2}}s_{\alpha_{2}}k\pm c_{\beta_{2}}s_{\alpha_{3}}\sqrt{1-k^{2}}=1.$ (23)
@@
  • $k=1\implies\alpha_{1}=\beta_{1}+2n\pi$. It is clear from Eqn. 23, that in this case we have $\cos\alpha_{3}\,\sin(\beta_{2}-\alpha_{2})=1$. This has two possible solutions: $\alpha_{3}=0$, $\alpha_{2}=\beta_{2}-\frac{\pi}{2}$ and $\alpha_{3}=\pi$, $\alpha_{2}=\beta_{2}-\frac{3\pi}{2}$. We can combine these conditions to read $\tan\alpha_{2}=-\cot\beta_{2}$.
@@
We have not studied the $k=-1$ case separately because there exists a clear mapping between the $k=1$ and $k=-1$ cases that does not affect any of the results in the current work. In the Regular hierarchy, changing $k$ from $+1$ to $-1$ is tantamount to the field redefintions $H_{1}\to H_{1}$, $H_{2}\to-H_{2}$, and $H_{3}\to-H_{3}$. Similarly in the Medial Hierarchy, this is equivalent to $H_{1}\to-H_{1}$, $H_{2}\to H_{2}$, and $H_{3}\to-H_{3}$, and in the Inverted Hierarchy case it is $H_{1}\to-H_{1}$, $H_{2}\to-H_{2}$, and $H_{3}\to H_{3}$. Thus, the $k=-1$ results can be gotten from the $k=1$ results by employing field redefinitions where the two Higgses other than the 125 GeV one change sign. Since the constraints have been arrived at by computing cross-sections of specific processes, all our results depend only on the square of the couplings and will not change. However, the $k=1$ and $-1$ cases can be potentially distinguished if there are more than one Feynamn diagrams contributing to a specific final state and thus, these two cases might contribute constructively or destructively thus altering the rates. However, we have checked that these considerations do not apply to any of the results in the present work.
@@ V Methodology
The Three Higgs Doublet Model (3HDM) presents a formidable challenge due to the curse of dimensionality, with its scalar potential governed by $15$ independent parameters. Our objective is to efficiently explore this $15$ dimensional parameter space to identify regions that satisfy theoretical constraints. Exhaustively searching for allowed points using a grid-based method would be computationally infeasible, potentially taking years. To circumvent this, we propose training a binary machine learning classifier  [^4] For a comprehensive review of ML techniques in particle physics, the reader is encouraged to consult [58, 59]. to predict whether a given point adheres to the constraints. However, the initial dataset generation reveals a substantial class imbalance, where most points fall into the disallowed category, thus skewing the dataset. The success of the machine learning model hinges on both the quality and representativeness of the training data. Crucially, points near the decision boundary—where the classifier exhibits the greatest uncertainty—carry more informational value than others. To optimize the selection of such data points, we employ an Active Learning [60] strategy. Our Active Learning algorithm significantly enhances the efficiency of the neural network classifier by focusing on a minimal yet highly informative subset of points. This process unfolds as follows:
@@ VII Discussion and Conclusions
To understand the regions of parameter space that are allowed by all experimental and theoretical constraints, we studied three cases: Regular Hierarchy wherein the lightest of the three Higgses is the SM-like Higgs, Medial Hierarchy with the second lightest being the 125 GeV Higgs, and Inverted Hierarchy in which the heaviest of the three is identified as the SM-like Higgs. Imposing all theoretical and experimental and theoretical constraints even in this simplified set-up is still a daunting task given the multidimensional nature of the 3HDM parameter space. To mitigate this, we adopt a machine learning strategy by designing an active learning algorithm to effectively delineate the decision boundary between allowed and disallowed regions. This reduces the computational time significantly, while also maximizing accuracy.
We find, under this set-up, that this model - though highly constrained, still does have a lot of parameter space that can potentially be probed at the LHC and other future colliders. Both the Regular and Medial Hierarchy cases have heavy scalar and pseudoscalar Higgs bosons in the range of $100-600$ GeV, potentially in the right ball park to warrant a dedicated LHC study. Interestingly, the case of having at least one CP-even Higgs lighter than the SM-like 125 GeV Higgs is not ruled out, as our results for the Medial Hierarchy case demonstrate. However, the more radical scenario of having two Higgses lighter than the 125 GeV SM-like Higgs is not favored by electroweak precision data. Nevertheless, the 3HDM remains a very interesting possibility in its own right, both from the point of view of the theoretical formulation (that can take many avatars depending on the discrete symmetries and the choice of Yukawas) and the associated collider phenomenology.
@@ Acknowledgements.
A.K. acknowledges the support from the Director’s Fellowship at IIT Gandhinagar. A.K. gratefully acknowledges Kousik Loho, Rameswar Sahu, and Lalit Pathak for their insightful discussions and valuable inputs during the course of this work. The authors also thank Dipankar Das for pointing out a typographical error in the manuscript. A.S. thanks Anusandhan National Research Foundation (ANRF) for providing necessary financial support through the SERB-NPDF grant (Ref No: PDF/2023/002572). S.K.R. acknowledges the support from the Department of Atomic Energy (DAE), India, for the Regional Centre for Accelerator-based Particle Physics (RECAPP), Harish Chandra Research Institute.
@@ Appendix A Charged Higgs Mass Matrix Calculation
$V_{C}^{mass}\supset\begin{pmatrix}\phi_{1}^{-}&\phi_{2}^{-}&\phi_{3}^{-}\end{pmatrix}\mathcal{M}^{2}_{\phi^{\pm}}\begin{pmatrix}\phi_{1}^{+}\\ \phi_{2}^{+}\\ \phi_{3}^{+}\end{pmatrix}$
@@ A.1 Transformation
$V_{C}^{mass}=\begin{pmatrix}\phi_{1}^{-}&\phi_{2}^{-}&\phi_{3}^{-}\end{pmatrix}\mathcal{M}^{2}_{\phi^{\pm}}\begin{pmatrix}\phi_{1}^{+}\\ \phi_{2}^{+}\\ \phi_{3}^{+}\end{pmatrix}$